a 75% salt solution is to be mixed with a 50% salt solution to get 12 liters of a 60% salt solution. how many liters of the 75% solution is needed.
Answer by lwsshak3(11628)

a 75% salt solution is to be mixed with a 50% salt solution to get 12 liters of a 60% salt solution. how many liters of the 75% solution is needed.
**
let x=liters of 75% solution to be mixed
12-x=liters of 50% solution to be mixed
..
75%x+50%(12-x)=60%*12
.75x+6-.5x=7.2
.25x=1.2
x=4.8
12-x=7.2
liters of 75% solution to be mixed=4.8
liters of 50% solution to be mixed=7.2